OpenStudy (anonymous):

help? 2 questions medal c:!

OpenStudy (anonymous):

1.A 2000 W toaster operates on a household circuit of 200 V. What is the resistance of the wire the make up the heating element of the toaster?

OpenStudy (ksaimouli):

Resistance = Voltage / Current

OpenStudy (anonymous):

0.1?

OpenStudy (ksaimouli):

Si unit of current is amp but they gave us power

OpenStudy (anonymous):

electric power (watts)=votlage*amperage, amperage=voltage/resistance so 2000W=200V*Amperage (10amp). 10amp=200V/resistance (20)

OpenStudy (anonymous):

20 ohms

OpenStudy (ksaimouli):

\[power=\frac{ V }{ I }\]

OpenStudy (ksaimouli):

solve for V

OpenStudy (ksaimouli):

from above equation

OpenStudy (anonymous):

ty :) @ksaimouli and @SalvadorV kk hold on

OpenStudy (anonymous):

you have the intensity formula wrong

OpenStudy (anonymous):

me?

OpenStudy (anonymous):

sorry the power formula

OpenStudy (anonymous):

P=V*I

OpenStudy (anonymous):

oh so I got it wrong?

OpenStudy (anonymous):

yes, Watts=Volts*Amps Amps=Volts/Ohms

OpenStudy (anonymous):

oh wait im confused .-. sorry

OpenStudy (anonymous):

watts=volts^2/ohms 2000=200^2/ohms ohms=200^2/2000 = 40 000/2000=20 ohms

OpenStudy (anonymous):

don't worry, we all make mistakes

OpenStudy (anonymous):

I feel stupid :l but what is ohms?

OpenStudy (ksaimouli):

resistance (unit)

OpenStudy (anonymous):

ohms is the measure for resistance

OpenStudy (anonymous):

oohh ok! ty for helping me c: do you mind helping with 1 more?

OpenStudy (anonymous):

and ksaimouli has his formula wrong .-.

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

that's why I got confused :c and kk

OpenStudy (anonymous):

2.An electronic device is rated at 0.35 W when connected to 150 V. What is the resistance of this device?

OpenStudy (anonymous):

see, its the same kind of question so \[0.35W=150V*amps\]

OpenStudy (anonymous):

so \[amps=\frac{ 0.35W }{ 150V }\]

OpenStudy (anonymous):

I was just about to put that lol

OpenStudy (anonymous):

ty again :)

OpenStudy (anonymous):

and we know that \[Amps=\frac{ volts }{ ohms }, 0.00233333333=\frac{ 150V }{ Ohms }\]

OpenStudy (anonymous):

remember ohms are the unit for resistance, normally noted with the sign\[\Omega \]

OpenStudy (anonymous):

here we have to substitute for ohms, so we pass them to the left side multiplying, and then we pass the 0.00233333333 dividing the voltage

OpenStudy (anonymous):

so we get \[\Omega=\frac{ 150V }{ 0.00233333333Amp }\]

OpenStudy (anonymous):

got it ! :) so =64285.72

OpenStudy (anonymous):

yea.

OpenStudy (anonymous):

sorry they left u so soon.

OpenStudy (anonymous):

um ok

OpenStudy (anonymous):

wow. nice pic.

OpenStudy (anonymous):

perfect! :D

OpenStudy (anonymous):

wow dude. ur late.

OpenStudy (anonymous):

.-. ok, whatever you say

OpenStudy (anonymous):

I'm actually trying to help people here instead of flirting you know?

OpenStudy (anonymous):

i'm helping finish the work u left unfinished. plus i could get her before u do anyway.

OpenStudy (anonymous):

I don't care about that, I have a girlfriend, plus just hopping in and saying "yes" is not exactly helping much

OpenStudy (anonymous):

i know. but i also know not 2 leave a girl hanging.